# MagicDistillation: Weak-to-Strong Video Distillation for Large-Scale Portrait Few-Step Synthesis

### v1

@@ Appendix E Pseudo Code of W2SVD
Appendix F Training Collapse of Vanilla DMD
Figure 8: Comparison between vanilla DMD and W2SVD in terms of training stability. It is worth noting that by the 400th iteration, vanilla DMD exhibits a noticeable change in the background, which serves as an early warning sign of impending training collapse. In contrast, W2SVD (w/ reg. loss) demonstrates enhanced stability, effectively mitigating such issues throughout the training process.
Here, we employ visualization to demonstrate the training collapse issue encountered with vanilla DMD. It is important to highlight that the number of iterations required to distill HunyuanVideo from 28 steps to 4 steps is remarkably small, demanding only a few hundred to a few thousand iterations with a batch size of 4. However, vanilla DMD suffers from training collapse after approximately 300 iterations, leading to the synthesis of uncontrollable videos. In Fig. 8, we present the videos synthesized by vanilla DMD in the 4-step scenario across different numbers of iterations, where this collapse phenomenon is clearly observable. In contrast, W2SVD effectively addresses this issue by aligning the real data distribution with the fake data distribution, while also leveraging ground truth supervision to entirely avoid this problem. As shown in In Fig. 8, W2SVD consistently produces high-quality portrait videos throughout the entire training process.
Appendix G Our Customized VBench
@@
Appendix H Additional Visualization
Figure 10: Visualization of W2SVD at the 4-step scenario. The leftmost frame of the video (i.e., the first frame) is the reference image. It can be seen that all videos have high quality and the 4-step generator does not synthesize uncontrollable videos.
Figure 11: Visualization of W2SVD at the 1-step scenario. The leftmost frame of the video (i.e., the first frame) is the reference image. It can be seen that all videos have relatively high quality and the 1-step generator does not synthesize uncontrollable videos.
Here, we showcase the video frames from the synthesized videos of I2V-HunyuanVideo distilled using our proposed W2SVD. The complete videos can be accessed on the project page. Specifically, we present video frames from two categories of synthesized videos: the 1-step and 4-step scenarios. For the 1-step scenario, the visualization results are provided in Fig. 11. Similarly, for the 4-step scenario, the visualization results are displayed in Fig. 10.

### v2

@@ 17       $\mathcal{L}_{\text{dis}}\leftarrow\text{discriminatorLoss}(D_{\xi},G_{\phi},\epsilon)$ // Eq 4
Algorithm 1 MagicDistillation Training Procedure
Appendix G Training Collapse of Vanilla DMD
Figure 9: Comparison between vanilla DMD and MagicDistillation in terms of training stability. It is worth noting that by the 400th iteration, vanilla DMD exhibits a noticeable change in the background, which serves as an early warning sign of impending training collapse. In contrast, MagicDistillation (w/ reg. loss) demonstrates enhanced stability, effectively mitigating such issues throughout the training process.
Here, we employ visualization to demonstrate the training collapse issue encountered with vanilla DMD. It is important to highlight that the number of iterations required to distill HunyuanVideo from 28 steps to 4 steps is remarkably small, demanding only a few hundred to a few thousand iterations with a batch size of 4. However, vanilla DMD suffers from training collapse after approximately 300 iterations, leading to the synthesis of uncontrollable videos. In Fig. 9, we present the videos synthesized by vanilla DMD in the 4-step scenario across different numbers of iterations, where this collapse phenomenon is clearly observable. In contrast, MagicDistillation effectively addresses this issue by aligning the real data distribution with the fake data distribution, while also leveraging ground truth supervision to entirely avoid this problem. As shown in In Fig. 9, MagicDistillation consistently produces high-quality portrait videos throughout the entire training process.
Appendix H Our Customized VBench
@@
Appendix I Additional Visualization
Figure 11: Visualization of various accelerated sampling methods under the 1-step scenario. The video quality and dynamic degree synthesized by MagicDistillation are visually superior to those produced by Euler, LCM, and vanilla DMD2.
Figure 12: Visualization of MagicDistillation at the 4-step scenario. The leftmost frame of the video (i.e., the first frame) is the reference image. It can be seen that all videos have high quality and the 4-step generator does not synthesize uncontrollable videos.
Figure 13: Visualization of MagicDistillation at the 1-step scenario. The leftmost frame of the video (i.e., the first frame) is the reference image. It can be seen that all videos have relatively high quality and the 1-step generator does not synthesize uncontrollable videos.
Here, we first present the video frames under the 1-step scenario and further showcase additional video frames from the synthesized videos of MagicDistillation. For the 1-step scenario, as shown in Fig.  11, the visual quality of MagicDistillation remains significantly better than that of the comparison methods, despite a slight decrease compared to the 4-step scenario. For more examples, the complete videos can be accessed on https://magicdistillation.github.io/MagicDistillation/. Specifically, we provide video frames from two categories of synthesized videos: the 1-step and 4-step scenarios. The visualization results for the 1-step scenario are presented in Fig. 13, while those for the 4-step scenario are displayed in Fig. 12.